# Directional and contra-directional coupling in Huygens’ metawaveguide microring resonators

### v2

@@ Introduction
Building on the previous bandgap calculations, we designed an apodized single-band filter with a peak bandwidth of = 10 nm. The length of the CDC was set to = 150 m, and a Gaussian apodization window was applied, with the grating edges reflecting 10 of the peak bandwidth ( 1 nm). The filter comprises three distinct sections, as shown in Fig. 5(a). In the central section of the device, corresponding to the apodized CDC, the mode travelling to the right through the Huygens’ waveguide is coupled backwards to the SWG waveguide. The backward coupling occurs within a 10-nm wide band centered at = 1550 nm. The SWG waveguide segments are modulated based on the data from Fig. 4(d) and following the target Gaussian bandgap profile. The results of these calculations are shown in Fig. 5(b). After coupling to the SWG waveguide, the mode travels to the left and transitions into a 450-nm wide strip waveguide over a length of = 10 m. The strip waveguide then bends away from the Huygens’ waveguide over a length of = 35 m and using a bending radius of R=20 m. Identical structures are also included at the output of the SWG waveguide to extract any directionally coupled light. Experimental results of the filter are shown in Fig. 5(c). Bandwidth remained fixed at 10 nm after fabrication, and the power loss was around 1 dB when referenced to the straight Huygens’ waveguide transmittance. We attribute the observed blue-shift of the center wavelength to deviations in the layout feature sizes during fabrication. An interference pattern was observed at the add port, which results from residual directional coupling between the Huygens’ and the SWG waveguides. However, this interference had minimal impact on the transmittance at the through port.
@@
To conclude this section, we experimentally assessed the behaviour of the designed apodized CDC when integrated with a racetrack resonator. The device under test is shown in the SEM images of Fig. 6(a). The coupling length was varied from 20 m to 300 m in several steps, which in turn modified the free spectral range (FSR) of the resonances from 2.3 to 0.5 nm, respectively. A more detailed view of the racetrack resonances is provided in Fig. S5 in Supporting Information Section V. The transmission spectra of the CDC-loaded racetrack resonators at the through, add and drop ports are presented in Fig. 6(b-g). The observed resonances exhibited an exceptionally low FSR, while the CDC also limited the light coupling spectrum of the racetrack to a 10-nm rejection band centered within the Huygens’ operation bandwidth. This demonstrated the ability to tailor the Huygens’ racetrack resonances with using an envelope function provided by an apodized SWG-assisted CDC.
@@ Conclusion
In summary, this work has introduced and experimentally validated a new class of integrated photonic components based on Huygens’-resonant microring resonators, and directional and contra-directional couplers, all engineered for operation at the 1550 nm telecommunication wavelength. By exploiting the unique dispersion and group velocity properties of Huygens’ waveguides, we demonstrated efficient evanescent coupling in micro-ring resonators implemented in both directional and contra-directional configurations, relevant for compact, high-performance add-drop filters. We also explored the broader implications of these structures for controlling group index and group velocity dispersion, pointing to promising applications in nonlinear and quantum photonics. A key innovation is the development of a hybrid subwavelength grating–Huygens’ contra-directional coupler, enabling backward coupling across resonant and non-resonant metawaveguides and supporting broad spectral rejection bandwidths. This advancement facilitates the design of FSR-free racetrack resonators, opening new pathways for spectral engineering. Altogether, we anticipate that this work will open new research avenues, enabling incorporation of Huygens’ waveguide in the future integrated photonic platforms for both classical and quantum optical communication, computing and sensing applications.
@@ Acknowledgements
The project was funded by the (1) National Research Council Canada (NRC)’s (i) Postdoctoral Fellowship Program, (ii) Collaborative Science, Technology and Innovation Program (CSTIP) Small Teams (ST-R2-01-02), and by the (2) Deutsche Forschungsgemeinschaft (DFG, German Research Foundation, project number 448835038).
M. Saad Bin-Alam acknowledges the support from the NRC’s Postdoctoral Fellowship in Metamaterial Integrated Photonics, and thanks Dr. Shurui Wang and Martin Vachon for their help in the experimental characterizations.

### v3

@@ Introduction
Building on the previous bandgap calculations, we designed an apodized single-band filter with a peak bandwidth of = 10 nm. The initial length of the CDC was set to = 150 m, which for a bandgap of 10 nm provides a calculated rejection of 30 dB, and a Gaussian apodization window was applied, with the grating edges reflecting 10 of the peak bandwidth ( 1 nm). Assuming a group index of ng = 4, the design targets a theoretical rejection ratio exceeding 60 dB to ensure efficient contra-directional power transfer to the drop port. The Gaussian apodization of the grating is employed to achieve low insertion loss and reduced out-of-band ripples. The filter comprises three distinct sections, as shown in Figure 5(a). In the central section of the device, corresponding to the apodized CDC, the mode traveling to the right through the Huygens’ waveguide is coupled backward to the SWG waveguide. The backward coupling occurs within a 10-nm wide band centered at = 1550 nm. The SWG waveguide segments are modulated based on the data shown in Figure 4(d) and following the target Gaussian bandgap profile. The results of these calculations are shown in Figure 5(b). After coupling to the SWG waveguide, the mode transitions into a 450-nm-wide strip waveguide over a length of = 10 m, , 46 periods of the SWG waveguide. The size of the SWG inner and outer segments is continuously adjusted to close the bandgap as needed by the Gaussian apodization. At the junction with the SWG–wire coupler, the SWG waveguide has the same inner and outer segment widths. The strip waveguide then bends away from the Huygens’ waveguide over a length of = 35 m and using a bending radius of R=20 m. Identical structures are also included at the output of the SWG waveguide to extract any directionally coupled light. The experimental results of the filter are shown in Figure 5(c). The bandwidth remained fixed at 10 nm after fabrication, and the power loss was around 1 dB when referenced to the straight Huygens’ waveguide transmittance. We attribute the observed blue-shift of the center wavelength to deviations in the layout feature sizes during fabrication. An interference pattern was observed at the add port, which results from residual directional coupling between the Huygens’ and the SWG waveguides. However, this interference had minimal impact on the transmittance at the through port.
@@
To evaluate the influence of the CDC on racetrack-resonator performance, devices with different CDC lengths were fabricated and characterized experimentally. A device under test is shown in the SEM images of Figure 6(a). The coupling length was varied from 20 m to 300 m in several steps, which in turn modified the of the resonances from 2.3 to 0.5 nm, respectively. A more detailed view of the racetrack resonances is provided in Figure S7 in Supporting Information Section VII. The transmission spectra of the CDC-loaded racetrack resonators at the through, add and drop ports are presented in Figure 6(b-g). The observed resonances exhibited an exceptionally low , while the CDC also limited the light coupling spectrum of the racetrack to a 10-nm rejection band centered within the Huygens’ operation bandwidth. This demonstrated the ability to tailor the Huygens’ racetrack resonances using an envelope function provided by an apodized SWG-assisted CDC. The power conservation in the CDC and the CDC-loaded racetrack resonators is discussed in Supporting Information Section VIII.
@@ Conclusion
In summary, this work has introduced and experimentally validated a new class of integrated photonic components based on Huygens’-resonant microring resonators, and directional and contra-directional couplers, all engineered for operation at the S- and C-band telecommunication wavelengths. By exploiting the unique dispersion and group velocity properties of Huygens’ waveguides, we demonstrated efficient evanescent coupling in micro-ring resonators implemented in both directional and contra-directional configurations, relevant for compact, high-performance add-drop filters. We also explored the broader implications of these structures for controlling group index and group velocity dispersion, pointing to promising applications in nonlinear and quantum photonics. A key innovation is the development of a hybrid subwavelength grating–Huygens’ contra-directional coupler, enabling backward coupling across resonant and non-resonant metawaveguides and supporting broad spectral rejection bandwidths. This advancement facilitates the design of FSR-free racetrack resonators, opening new pathways for spectral engineering.
Our Huygens’ metawaveguide devices differ from conventional photonic-crystal (PhC) waveguides in that light guidance and dispersion arise from locally resonant dielectric antennas rather than from extended Bloch modes supported by a large periodic lattice. In PhCs, phenomena such as slow light, negative group velocity, and contra-directional coupling depend on long-range Bragg interference across many unit cells, making them sensitive to fabrication disorder and difficult to tune after fabrication. In contrast, Huygens metawaveguides employ directionally scattering electric–magnetic resonators whose near-field coupling can be engineered locally. This opens a path towards a potentially reduced device footprint, flexible dispersion control, improved tolerance to fabrication imperfections, and dynamic reconfigurability, making Huygens architectures more versatile than traditional PhC-based waveguide structures.
The Huygens’ metawaveguide devices operate over a bandwidth approaching 100 nm, which is fundamentally limited by the spectral overlap of the ED and MD resonance linewidth. In future work, this bandwidth could be further extended by engineering the antenna geometry to simultaneously excite higher-order electric and magnetic quadrupole resonances (EQ and MQ) within a comparable wavelength range, thereby broadening the effective Huygens regime. Additional bandwidth-limiting phenomena arise from the photonic bandgap at low frequencies and the onset of diffraction losses at high frequencies near the light line, which can be optimized through judicious design. In addition to bandwidth enhancement, further improvements in add–drop filter performance could possibly be realized by increasing the radii of the resonators’ curved sections to mitigate bend-induced scattering losses and enhance overall extinction ratio and transmission efficiency.
Altogether, we anticipate that this work will open new research avenues, enabling incorporation of Huygens’ waveguide in the future integrated photonic platforms for both classical and quantum optical communication, computing and sensing applications.
Funding Sources
The project was funded by the (1) National Research Council Canada (NRC)’s (i) Postdoctoral Fellowship Program, (ii) Collaborative Science, Technology and Innovation Program (CSTIP) Small Teams (ST-R2-01-02), and by the (2) Deutsche Forschungsgemeinschaft (DFG, German Research Foundation, project number 448835038). M. Saad Bin-Alam acknowledges the support from the NRC’s Postdoctoral Fellowship in Metamaterial Integrated Photonics.
@@ Acknowledgements
The authors thank Dr. Shurui Wang and Martin Vachon for their help in the experimental characterizations.
@@
Associated Content
A pre-print version of this manuscript is available at arXiv:2510.15845.
Bin-Alam, M. S., Sirmaci, Y. D., Fernández-Hinestrosa, A., Zhang, J., Dolgaleva, K., Boyd, R. W., Luque-González, J. M., Pertsch, T., Staude, I., Schmid, J. H., Cheben, P., Directional and contra-directional coupling in Huygens’ metawaveguide microring resonators. 2025, arXiv preprint arXiv:2510.15845. https://doi.org/10.48550/arXiv.2510.15845 (March 27, 2026).